OpenStudy (adamk):

From what height above the surface of the earth should an object be dropped to initially experience an acceleration of 0.54g?

OpenStudy (michele_laino):

I think that we can apply this formula: \[\Large g' = K\frac{M}{{{{\left( {R + h} \right)}^2}}}\] where \(g'=0.54\times g\) \(M\) is the mass of earth, and \(R\) is the radius of earth. \(K\) is the constant of Newton, and \(h\) is the requested height

OpenStudy (adamk):

Ahhh that formula. Thank you.